### v3

# Evidence of Scaling Regimes in the Hopfield Dynamics of Whole Brain Model

Abstract
It is shown that a Hopfield recurrent neural network exhibits a scaling regime, whose specific exponents depend on the number of parcels used and the decay length of the coupling strength. This scaling regime recovers the picture introduced by Deco et al., according to which the process of information transfer within the human brain shows spatially correlated patterns qualitatively similar to those displayed by turbulent flows, although with a more singular exponent, $1/2$ instead of $2/3$. Both models employ a coupling strength which decays exponentially with the Euclidean distance between the nodes, informed by experimentally derived brain topology. Nevertheless, their mathematical nature is very different, Hopf oscillators versus a Hopfield neural network, respectively. Hence, their convergence for the same data parameters, suggests an intriguing robustness of the scaling picture. Furthermore, the present analysis shows that the Hopfield model brain remains functional by removing links above about five decay lengths, corresponding to about one sixth of the size of the global brain. This suggests that, in terms of connectivity decay length, the Hopfield brain functions in a sort of intermediate “turbulent liquid”-like state, whose essential connections are the intermediate ones between the connectivity decay length and the global brain size. The evident sensitivity of the scaling exponent to the value of the decay length, as well as to the number of brain parcels employed, leads us to take with great caution any quantitative assessment regarding the specific nature of the scaling regime.
keywords:
Hopfield Network , Whole-Brain Model , Neural Mass Model , Scaling Regimes
journal: Neural Networks
[…]
[a]organization=Center for Life Nano- and Neuro-Science, Istituto Italiano di Tecnologia,addressline=Viale Regina Elena 201, city=Rome, postcode=I-00161, country=Italy
[…]
[b]organization=Istituto di Scienze del Patrimonio Culturale, Consiglio Nazionale delle Ricerche,addressline=Strada della Neve s.n.c., Via Salaria km 29.300, city=Montelibretti (RM), postcode=I-00010, country=Italy
[…]
[c]organization=Istituto per le Applicazioni del Calcolo del Consiglio Nazionale delle Ricerche,addressline=via dei Taurini 19, city=Roma, postcode=I-00185, country=Italy
[…]
[d]organization=Dipartimento di Fisica, Universitá di Roma “La Sapienza”,addressline=P.le Aldo Moro 5, city=Roma, postcode=I-00185, country=Italy
{highlights}
The mesoscale Hopfield model of the global brain collective dynamics presents a scaling regime.
It recovers the scaling picture recently introduced by Deco et al. 2020 based on a Hopf oscillator model. The near-quantitative convergence of the two models points to an intriguing robustness of the scaling picture.
A scaling regime implies that the brain functions in a sort of intermediate “turbulent liquid”-like state.
It also shows that the scaling exponent associated with such collective patterns is highly sensitive to the decaying length of the brain connectivity, as well as to the number of brain parcels.
The global brain Hopfield model is shown to remain functional upon removing connections above five decay length scales.
1 Introduction
In a recent paper, Deco and Kringelbach (2020) argued that the information transfer within the human brain may proceed in close analogy with the mechanisms which govern mass and energy transport in turbulent fluids. The analogy refers to the emergence of correlated spacetime patterns which would facilitate and possibly optimize information transfer across the brain, in the “same” way as fluid turbulence enhances energy transport across fluids. This is interesting in many respects, not least the prospects for future neurocomputers (Schuman et al., 2022).
Based on an extended dynamical system of oscillators (Hopf model), supplemented by large-scale neuro-imaging empirical data on the connections between different brain areas, Deco and Kringelbach (2020) numerically showed that the statistical correlations of human brain signals may exhibit a scaling regime similar to the one observed in turbulent fluids.
At a closer scrutiny, though, the analogy does not appear to be quantitative, meaning that the spatial scaling exponents of the second order structure function $S_{2}(d)$=$\langle(O(d)-O(0))^{2}\rangle$$\sim d^{\alpha}$, $O$ being a suitable order parameter, are not the same, about $\alpha$$\sim$1/2 for the Hopf model of the human brain against $\alpha$=2/3 for fluid turbulence (Benzi et al., 1993). Hence “brain turbulence” appears to be more singular than fluid turbulence.
Yet, a transient scaling regime in space is found in either cases, which is intriguing per se, regardless of the specific value of $\alpha$.
In this paper, we examine the robustness of the aforementioned analogy by conducting a similar analysis based on the same neuroimaging data for brain connectivity, but using a very different model: a Hopfield neural network. Additionally, we measure the scaling exponents as a function of the connectivity decay length ($\delta$), discovering a much richer scenario, including a dependence on the number of parcels (network nodes) ($N$), to be detailed later in the paper.
The recurrent Hopfield neural network model is a well-established model for small-scale neuronal networks (Amari, 1972; Little, 1974; Hopfield, 1982) and a particularly minimal discrete-time recurrent neural network model (Grossberg, 1967; Rumelhart and McClelland, 1986). It is a neural network model based on binary McCulloch-Pitts neurons (Amit et al., 1985; Brunel, 2016; Hillar and Tran, 2018; Hillar et al., 2021) in which every processing unit ($i$) is connected to all the other ones ($j$) through a set of weights ($J_{ij}$). Most of the literature considers fully connected and symmetric Hopfield networks since this simplifies the mathematical analysis of the model, although recently the investigation of diluted networks (Stauffer et al., 2003; Brunel, 2016; Kim et al., 2017; Folli et al., 2018; Leonetti et al., 2020) and asymmetric networks (Gopalsamy and zhong He, 1994; Xu et al., 1996; Chen and Amari, 2001; Franca and Yang, 2000; Zheng et al., 2010; Szedlak et al., 2014; Folli et al., 2018; Leonetti et al., 2020) was introduced.
Usually, a recurrent Hopfield neural network is trained to store patterns or memories as steady-state attractors with the Hebbian prescription. In this case, the coupling matrix $J_{ij}$ takes the form of the sum of dyadics, and therefore it is symmetric and fully connected. The maximum storage capacity of such a network is found to scale linearly with the number of nodes $N$. The capacity can be increased by the process of learning (Hebbian learning) whereby the simultaneous activation of two nodes reinforces their connection. This is inspired by biological neural networks where the simultaneous activation of neurons leads to increments in synaptic strength (Hebb, 1949), and it allows the network to store a larger number of patterns as steady-states (Amit et al., 1985). Yet, this quantity ($C$) remains linear with $N$: $C\sim N$, being the proportionality factor of the order of unity. Lately, researchers have also investigated how unlearning methods can improve the stability of the stored patterns Fachechi et al. (2019); Benedetti et al. (2022); Benedetti and Ventura (2023); Agliari et al. (2019, 2024). It is worth emphasizing that a vast literature, tracing back to the seminal papers by Gardner (1986) and Hopfield et al. (1983), treat the stationary points of the Hopfield dynamics as “memories”, but in the context of whole brain dynamics, and in our work, we will not consider this interpretation.
Interestingly, it has been recently demonstrated both numerically (Folli et al., 2018; Gosti et al., 2019) and theoretically (Hwang et al., 2019, 2020) that random, non-Hebbian, coupling matrices have a number of steady-states much larger than the Hebbian ones. Indeed the number of steady-states $C$ of random matrices scales exponentially with $N$: $C\sim e^{\Sigma N}$ being $\Sigma$ the complexity. The latter quantity is found to be maximized for asymmetric and diluted matrices, with values that match the connectivity found in the hippocampus areas.
The approach outlined by Deco and Kringelbach (2020) diverges from both the Hebbian and random connectivity models. It relies on a coupling matrix, which is derived from experimental observations of the connections between human brain regions. This approach assumes that the coupling strengths are all positive, deterministic, and specifically proportional to a given function of the distances between nodes. $J_{ij}=e^{-d_{ij}/\delta}$, $d_{ij}$ being the Euclidean distance between nodes $i$ and $j$ in the network, and $\delta$ a decay length, which, in Deco and Kringelbach (2020) are kept fix to match the empirical HCP dMRI tractography of the human brain: $\delta=5.55$ $\mathrm{m}\mathrm{m}$.
At variance with the latter approach, where the dynamics is based on the Hopf oscillator model, here we use the same coupling matrix of Deco and Kringelbach (2020), but employ the Hopfield dynamics.
The use of Hopfield networks to investigate the collective dynamics of the brain activity is not new (Trappenberg, 2022). Indeed, Deco et al. (2012), and Golos et al. (2015) used Hopfield models (or Ising-Spin models with Glauber dynamics) as a neural mass model to study the complex landscape of brain activation dynamics. Neural mass models schemes do not interpret the nodes of the Hopfield network as single spiking neurons but rather as aggregates of neural mass (parcel), namely mesoscale brain regions switching between active and rest states. In the active states, many neurons simultaneously fire while in the rest states most neurons are silent. These models do not aim at establishing a rigorous connection between the microscopic description of the brain and a corresponding mesoscopic one, but posit instead a physically-informed heuristic picture of the collective brain dynamics.
Broadly speaking, the Hopf model is based on the idea of reproducing the empirically observed information transfer between areas that takes place through (partial) synchronization of collective oscillations. In this framework, the natural collective variables describing the global brain dynamics are mesoscale oscillators obeying the Hopf dynamics. The Hopfield model used in this work draws inspiration from a very different idea, namely to carry over the discontinuous divide between active and rest states from the microscale level of single neurons to the mesoscale level of neural parcels, as discussed above. Formally, this amounts to assuming that coarse-graining of the Hopfield model would leave the model invariant but with a different set of weights, in our case an exponential decay. Such model proves useful to show that the dynamical repertoire of brain activity composed of sudden avalanches and bursts, can be described in terms of critical systems perturbed by noise (Deco et al., 2012; Golos et al., 2015).
Due to the significant difference between the fundamental assumptions behind these two models, inspecting their commonalities, as well as their points of departure, starting from the same empirical connectivity information, is of decided interest to improve the mesoscale description of the global brain dynamics,
Our main conclusions are summarized as follows:
[…]
ii) At the $\delta$ value investigated in Deco and Kringelbach (2020) ($\sim 5.5$ mm) and for $N=1000$, we found $\alpha\sim 2/5$, close to the one reported by Deco and Kringelbach (2020) [^1] In Ref. Deco and Kringelbach (2020) the used parameter is the inverse of the decay scale, $\lambda=1/\delta$.;
iii) For $N=1000$ the scaling exponent $\alpha$ shows a sigmoid-like trend with the decay length $\delta$. Importantly, for $N=1000$, the functional relation $\alpha(\delta)$ shows a steep increase in the region around $\delta\sim 5.55$ mm, the value indicated in Deco and Kringelbach (2020) as the physiologically relevant one from the analysis of empirical HCP dMRI tractography with a $N=1000$ parcellation. This means that the specific value of the decay length has a significant impact on the scaling exponent; for instance, a slight increase to $\delta$=5.88 mm would yield a scaling exponent slightly above $2/3$, the value associated with homogenoeus incompressible turbulence. Conversely, a slight decrease to $\delta$=5 mm, would yield $\alpha$$\sim$0.1, a much less organized regime.
iv) the sigmoid dependence of $\alpha$ on $\delta$ changes with $N$, becoming steeper and narrower at increasing $N$.
[…]
2 Discrete-time Hopfield Recurrent Neural Networks
Following (Amari, 1972; Little, 1974; Hopfield, 1982; Amit et al., 1985; Folli et al., 2017; Gosti et al., 2019; Leonetti et al., 2020; Gosti et al., 2024), we consider a network of $N$-binary nodes that can be either firing or silent. The state of each node is a binary variable $s_{i}$, with $i=1,2,\ldots,N$, such that $s_{i}=-1$ or $s_{i}=1$, so that the network state is represented by a binary string $\mathbf{s}=(s_{0},s_{1},\ldots,s_{N})$. Neurons interact through the connectivity matrix $\mathbf{J}$, with matrix elements $J_{ij}$ following the same exponential dependence $J_{ij}=e^{-d_{ij}/\delta}$ used in (Deco and Kringelbach, 2020). We assume a discrete time $t$, and, at each time step $t$, the evolution of the neuron state $s_{i}(t)$ is given by the following non-linear dynamic equation,
$s_{i}(t+1)=sign\left[\sum_{j=1}^{N}J_{ij}s_{j}(t)\right],$ (1)
where the activation function $sign(x)$ is defined to be 1 for $x$$\geq$0, and -1 otherwise. Consequently, whenever the summation of the inputs on node $i$ is above zero, the node is active (“fires”), otherwise, it remains silent.
We determined the couplings matrix $J_{ij}$ using Schaefer’s cerebral cortical parcellation atlas as in (Deco and Kringelbach, 2020). This parcellation was obtained from the analysis of a large dataset of neural activity dynamics with a sample size of 1489 (Schaefer et al., 2018). This dataset is publicly available. We used Schaefer’s parcellations with $N$ ranging from $200$ to $1000$. The Schaefer’s parcel centroids are mapped onto the MNI152 volumetric space [^2] Github repository dataset url Schaefer et al. (2018): https://github.com/ThomasYeoLab/CBIG/tree/master/stable_projects/brain_parcellation/Schaefer2018_LocalGlobal/Parcellations/MNI/Centroid_coordinates. The code for reproducing the simulations can be found in the git hub repository https://github.com/ggosti/HopBrain.
3 The Structure Factor $S_{2}(d)$
We used Deco and Kringelbach (2020) adaptation of Kolmogorov’s concept of structure functions. In turbulence, the Kolmogorov’s structure factor applies to the longitudinal velocity but in our case it represents the activity of the parcel as measured through the BOLD signal in a fMRI experiment.
To compute the structure factor we use the auxiliary covariance function $B(d)$ defined as:
$B(d)=\sum_{ij}\frac{1}{{\cal{N}}(d)}\sum_{d_{ij}\in d}s_{i}s_{j}$ (2)
where the function ${\cal{N}}(d)$ represents the number of $i,j$ pairs at distance $d$. The notation $d_{ij}\in d$ indicates that the distance between the nodes must be “close enough” to $d$. To this aim, it should be kept in mind that in Schaefer’s cerebral cortical parcellation atlas (Schaefer et al., 2018), $d$ was recorded with 2 mm precision, thus there are multiple $i,j$ pairs with the same $d$. Therefore we can bin the existing nodes’ distance and map them to the natural numbers, so as to use of the Kronecker delta $\Delta(ij|d)$, which takes value $1$ if $d_{ij}$ is in the bin identified by $d$, and is 00 otherwise.
As a result, we have:
$B(d)=\frac{1}{{\cal{N}}(d)}\sum_{ij}s_{i}s_{j}\Delta(ij|d)$ (3)
and
${\cal{N}}(d)=\sum_{ij}\Delta(ij|d)$ (4)
The function $B(d)$ relates to the structure factor as follows (Deco and Kringelbach, 2020):
$S_{2}(d)$ $=$ $2\left[B(0)-B(d)\right]$ (5)
The Hopfield dynamics is deterministic, and since the phase space of the nodes’ states is finite, the asymptotic trajectories must necessarily be periodic. Therefore, as also discussed in (Folli et al., 2017), Hopfield Recurrent Neural Networks, given an initial state, consistently end up into a specific steady state (“attractor”). The steady states are deterministic attractors composed of either limit cycles formed by repeating sequence of states, or a single stationary state. Given that in our case the couplings $J_{ij}$ are symmetric, we expected theoretically, and experimentally confirmed that the steady states are always composed of a single fixed point. In the absence of a further external input (“external stimuli”), and given sufficient time, the network gets stuck in an attractor and the whole activity is constant in time, hence no time summation was taken in the previous equations. This is of course a very crude representation of the brain dynamics, which receives continuously “internal” and “external” inputs, that perturb the system to constantly explore a landscape shaped by the collection of steady states. Given this context, Deco et al. (2012), and Golos et al. (2015) show how the whole set of attractors allows us to reproduce the dynamical repertoire and the functional correlations of the brain activation dynamics.
Scaling exponents vs. the decay length
Figure 1: Structure factor $S_{2}(d)$ for $N=1000$ nodes and $N_{r}=1000$ realizations respectively at $\delta=5.55$ mm (A), and $\delta=5.88$ mm (B). The values $S_{2}(d)$ were binned over equal sized intervals. The dashed lines indicate different $\alpha$ values: $\alpha=2/3$ which corresponds to turbulence, $\alpha=1/2$, the value obtained in (Deco and Kringelbach, 2020), and $\alpha=2/5$. We estimate the scaling exponent $\alpha$ with a linear regression of $\log(S(d))\sim\alpha\log(d)+\beta$ in the interval range $[2.7-33.1]$ mm, as used in (Deco and Kringelbach, 2020), and we obtain $\alpha\approx 2/5$ for $\delta=5.55$ mm, and $\alpha\approx 2/3$ for $\delta=5.88$ mm.
Given different values of the decay length $\delta$, we have run the Hopfied model for a system composed by $N=1000$ parcels (with pairwise distances consistent with the Schaefer’s cerebral cortical parcellation atlas (Schaefer et al., 2018)). The structure functions have then been calculated by following the system until the fixed points were reached, and by averaging over $N_{r}$=1000 randomly chosen initial states. As examples, figure 1 shows the structure factor $S_{2}(d)$ for respectively $\delta$=5.55 mm (Fig. 1A) and $\delta$=5.88 mm (Fig. 1B). The black points with the error bars represent the average $S(d)$ computed over 1000 different realizations. The dashed green line corresponds to $\alpha$=2/3, the orange line corresponds to $\alpha$=1/2, and the blue line corresponds to $\alpha$=2/5. The $\alpha$=1/2 value (orange line) is the value estimated in (Deco and Kringelbach, 2020). The $\alpha$=2/3 is the exponent associated with a turbulence. Due to the significant level of statistical fluctuations, we have estimated the average exponent in two different ways, first as the correct ensemble average $\langle S_{2}(d)\rangle$$\sim d^{\alpha}$ and also as an arithmetic average of the single realization exponents, $\overline{\alpha}$=$N_{r}^{-1}\sum_{r=1}^{N_{r}}\alpha_{r}$. Even though the latter is not correct, the fact that both yield very similar values attests to the robustness of the average exponent, notwithstanding sizeable fluctuations. For consistency with (Deco and Kringelbach, 2020), we estimate the scaling exponent $\alpha$ fitting the points in figure 1 in the interval 2.7 - 33.1 mm and we obtained $\alpha\approx$2/5 for $\delta$=5.55 mm, and $\alpha\approx$2/3 for $\delta$=5.88 mm.
Figure 2: A) The blue line shows the average slope of $\alpha=\log S(d)/\log d$ for different $\delta$s, given the connectivity matrix $J_{ij}=\exp{(-d_{i,j}/\delta)}$. The horizontal dotted lines correspond to $\alpha_{RW}=1$ and $\alpha_{T}=2/3$, corresponding to random walk and homogeneous incompressible turbulence, respectively, while the dot-dashed horizontal line corresponds to Deco and Kringelbach (2020) “brain” value $\alpha_{D}=1/2$. The orange line shows the function $\alpha(\delta)$ with a connectivity matrix obtained by randomly shuffling all the weights $J_{ij}$, which results into a basically uncorrelated signal, indicating the importance of the metric structure of the weights. The plot stops at $\delta=10$ mm, a region where the signal is highly correlated ($\alpha\sim 1.2$), but still below the smooth regime marked by $\alpha=2$. Conversely, a brain below $4$ mm, would exhibit “chaotic” behaviour, with scaling exponents close to zero, i.e. no scaling regime. The bottom panels show the neuron activity mapped on the Schaefer’s cerebral cortical parcellation atlas (Schaefer et al., 2018) coordinates for (B) $\delta=4$ (“chaotic”), (C) $\delta=5.99$, and (D) $\delta=6.66$ (random walk), respectively.
We repeated the previously depicted procedure for a sequence of values of the decay length in the range $[3-10]$ mm and measured the associated scaling exponents. The values of $\alpha(d)$ for different realizations are shown in Fig. 2 as blue points. The code to reproduce the simulation and the 3D model files (gltf) are available at the repository https://github.com/ggosti/HopBrain. This code also shows how to generate 3D maps for the Schaefer parcellation’s nodes activation states inspection, and how to export gltf files that can be viewed in immersive environments using open source webXR apps such as Aton which are part of the H2IOSC open cloud (Fanini et al., 2021; Fanini and Gosti, 2024) (Fig. 2B-D). Figure 2B-C can be accessed with a web-browser respectively with the following urls: https://aton.ispc.cnr.it/s/ggosti/20240612-6zkx4gyvi for $\delta=4$ (“chaotic”), https://aton.ispc.cnr.it/s/ggosti/20240612-ebn1rod5p for $\delta=5.99$, and https://aton.ispc.cnr.it/s/ggosti/20240612-a8e3initf for $\delta=6.66$ (random walk).
Several comments are in order. First, we observe a clear increasing trend of $\alpha$ with $\delta$, meaning that small decay lengths, i.e. localised connectivity, promotes irregular patterns. For instance, below $\delta\sim 4$ mm, the scaling exponent is basically zero, corresponding to complete randomness. This was checked independently, by running simulations in which the couplings $J_{ij}$ were randomized though keeping the same distribution of the elements of the matrix $J$ (flat orange line at the bottom). We were able to simultaneously randomize the couplings $J_{ij}$, and keep the same value distribution by randomly shuffling the $J_{ij}$ values. At the opposite end, with $\delta$=10 mm, we measure $\alpha$$\sim$1.2, which is still well below the smoothness threshold $\alpha_{S}$=2, but above the random walk value $\alpha_{R}$=1, and much above the Kolmogorov turbulence value $\alpha_{T}$=2/3, which is in turn larger than $\alpha_{D}$$\sim$1/2, the value obtained by Deco and Kringelbach (2020) at $\delta$$\sim$5.55 mm. Importantly, in the intermediate regime, around the physiological value $\delta$=5.55 mm, the present Hopfield model delivers $\alpha_{H}$$\sim$2/5, different but still close to Deco et al. value. A visual extrapolation of Fig. 2A seems to indicate that the scaling regime would remain non-smooth ($\alpha<2$) even in the limit of an "infinite-brain" ($\delta\to\infty$) the global brain ($d\sim 25\delta$) being pretty close to the infinite-brain limit. Note however that for $\delta>10$ mm we start getting cases in which $B(d)\sim B(0)$, so that the numerical calculation of the exponent diverges and becomes less and less accurate (a fully correlated signal with $B(d)=B(0)$ yields formally $\alpha\to\infty$.)
Figure 3 shows the pair correlation for different values of $\delta$. For $\delta<5$ mm, the pair correlation shows a basically uncorrelated activity, except at short range. For $\delta=5.26$ mm, long-range pair correlations start to emerge and finally, for larger $\delta$, the long-range pair correlations become even stronger. In the limit of very large $\delta$, stationary states emerge in which the system freezes into a single fully-ordered state, with all nodes states either $+1$ or $-1$. As mentioned above in the fully-ordered state with no noise $\alpha$ is infinite.
Figure 3: Pair-correlation functions $B(d)$ for different $\delta$s. At small $\delta$ the signal is largely uncorrelated, while at increasing $\delta$, correlations start to emerge, although affected by significant amount of statistical fluctuations.
Our analysis confirms that there is indeed a intermediate scaling regime in the organization of the brain patterns which reminds fluid turbulence, yet with a higher degree of spikiness. Indeed, by sticking to the physiological value $\delta=5.55$ mm, the Hopfield brain appears more irregular (lower scaling exponent) than a random walk and also of homogeneous incompressible turbulence. The most important feature, though, is the sigmoidal dependence of the relation $\alpha(\delta)$ with the rampup region between about $4$ and $7$ mm, centered around $\delta\sim 5.9$. This is indeed a physiologically relevant scale of local regions in the brain. Furthermore, this rampup region unveils the major sensitivity of the scaling exponent to small changes of the decay length, indicating that small changes of the decay length lead to fairly different scaling regimes. In particular, for $\delta$ below $4$ mm, the activity appears to be nearly chaotic, whereas, for $\delta$ above $8$ mm, it becomes kind of “rigid”, i.e. globally correlated.
4 Thresholding the connections
Having assessed that delocalization (large $\delta$s) promotes a correlated response, it is natural to inquire about the role of short versus long-range couplings in promoting the correlated patterns sustaining the scaling regime.
To this purpose, we performed a series of simulations by progressively removing pairs of nodes whose coupling strengths lie below a given threshold, namely $J_{ij}<J_{th}$, where the coupling strengths vary in the range $[0,1]$. Since $J_{ij}$ decays exponentially with the pair distance a threshold $J_{th}$ excludes pairs beyond a distance
$d_{th}=\delta ln{(1/J_{th})},$ (6)
thereby retaining only a fraction $1-J_{th}$ of the full set of interacting nodes. Thus, to a certain dilution corresponds a certain cutoff distance above which all couples are disconnected. The result of this progressive dilution on the scaling exponent are reported in Figure 5. Here the dilution $\rho$ is defined as the fraction of disconnected nodes $i,j$, $J_{ij}=0$, and the result on the scaling exponent is measured comparing the undiluted exponent $\alpha(0)$ with the exponent $\alpha(\rho)$ at that dilution.
Figure 4: The average slope of $\log S(d)$ for different dilution levels $\alpha(\rho)$, normalized to its threshold-free value $\alpha(0)$ at that given $\delta$. The dilution is measured as the fraction of disconnected pairs $i,j$, $J_{ij}$ elements with value 0. The second $y$-axis on top shows the cutoff distance in mm for which at a certain dilution $\rho$ the pairs at a larger or equal distance are disconnected. The differently colored curves indicate simulations with different $\delta$ values. The graph shows that within the measurement error, the system is not affected by the dilution of the edges, up to the removal of more than the $95\%$ of the connections, in descending order of distance. This shows that, in view of the exponential connectivity $J_{ij}=e^{-d_{ij}/\delta}$, the onset of collective patterns is sustained mostly by the close connections, within about $4\delta$.
From this figure, a smooth trend at increasing $\rho$ is observed, with the scaling exponent starting to display a significant decrease around $\rho=\left[0.95,0.98\right]$, when just 2 or 5 percent of the nodes are left, corresponding to $d_{ij}>4.6\delta$. With $J_{th}=0.10$, corresponding to $d_{ij}>-\delta ln(0.1)\sim 0.46\delta$, the scaling exponent is basically halved, indicating a major loss of correlation. This is expected since $J_{th}=0.10$ cuts out all but shortest range interactions. The conclusion of this dilution analysis is that the brain response is carried almost entirely by interactions up to about five decay lengths, namely about $1/6$ of the size of the global brain. This sounds reasonable, as it strikes a plausible compromise between short-range ($d<4\delta$) and long-range ($d>6\delta$) interactions.
It is interesting to observe that a “solid” (strongly correlated) brain, such as the one that results for $\delta>10$ mm, and low dilution $\delta$, would be too homogeneous to be able to store and process the amount of information required to function properly. On the opposite side, for too short connectivity length, the brain would behave like a “gas” of disconnected nodes, hence incapable of collective behavior, which is key to its proper functioning. It appears like the physiological length is achieving an optimal compromise between these two opposite, order-disorder, trends. With a daring but conducive metaphor, we could speculate that the brain works in a sort of turbulent liquid-like state, although more spiky than actual turbulent fluids.
5 Changing The Number of Parcels
To assess the impact of the parcellation size on the scaling picture discussed in this paper, we gathered Schaefer 2018 Parcellations (Schaefer et al., 2018) with different sizes $N$ and we run numerical simulations with $N_{r}=1000$ replicas and different values of $\delta$. This allowed us to inspect the relation between $\alpha$ and $\delta$ for different values of the size $N$. The result is the relation reported in Fig. 5, which shows a progressive regression of the transition regions towards smaller values of $\delta$, combined with a corresponding steepening of the transition.
Figure 5: The average slope of $\alpha$ as a function of $\delta$ for Schaefer 2018 Parcellations with different sizes $N$ (Schaefer et al., 2018). Given each size $N$, we fitted the points with the sigmoid functions in Eq. (7).
To provide a quantitative measure of this process, we fitted the points at given values of $N$ with the sigmoid function:
$\alpha(\delta)=\frac{\alpha_{\infty}}{1+\exp{(-k(\delta-\delta_{o}))}}.$ (7)
where $\alpha_{\infty}$ represents the plateau value of the scaling exponent at infinite decay length, $\delta_{0}$ is the transition scale and $k$ is the steepness of the transition region.
The sigmoid function captures the functional relation of the points reasonably well even though, as we discussed earlier on, the values $\alpha(\delta)$ do not fully converge in the top branch because for $\delta>10$ mm fully aligned configurations prevail, with correspondingly divergent values of $\alpha$. Thus, we fitted the function in three steps: first, we estimated $\alpha_{\infty}$ on the plateau points for $N=800,900,1000$. Then, we obtained $\delta_{0}$ and $k$ by fitting the sigmoid function $\alpha(\delta)$ on all points using a nonlinear least squares fit. We finally obtained $\delta_{0}$ and $k$ as a function of $N$ via a linear least square fit in loglog scale, and found $\delta_{0}\sim N^{-0.379}$ and $k\sim N^{0.328}$, with a $R^{2}$-value $0.999$ and $0.925$, respectively.
Based on the above scaling of the parameters, one would conclude that the sigmoid tends to a sign function centered in zero in the thermodynamic limit $N\to\infty$. In plain words, the transition length goes to zero and the transition region becomes infinitely steep. Hence, in the infinite $N$ thermodynamic limit, the system is fully correlated (“rigid”) even for vanishing decay lengths, an extreme instance of long-range ordering.
But the brain is not infinite, hence it is plausible to expect that there should exist optimal parcellation size $N_{opt}$ providing a correspondingly optimal Hopfield mass model approximation to the empirical correlations provided by a realistic model of structural connectivity. Due to its inherently coarse-grained nature, above $N_{opt}$, the Hopfield network whole brain mass model is no longer expected to provide a trustworthy approximation of the collective brain dynamics.
Even if we cannot pin down such an optimal value, we expect it to lie between $N=10^{3}$, the largest parcellation presently available to us and $N<10^{6}$, which provides a plausible upper bound to the number of functioning “modules” in the human brain (number of neurons, $10^{11}$, divided by the typical module size, $10^{5}$ neurons).
Given the actual scalings provided by our fit, a three orders of magnitude increase in the parcellation size would lead to about ten times smaller values of both $\delta_{0}$ and thickness $1/k$, thus placing the transition length $\delta_{0}$ at about $0.5$ mm, with a strong sensitivity of the scaling exponent confined within a region of lengthscale about $0.1$ mm wide. In other words, going to larger parcellations makes the scaling exponent increasingly sensitive to small changes of the decay length.
Future experimental developments will allow us to explore finer parcellations with larger $N$ values and test the predictions made in this work. Furthermore, from these and other independent considerations Kaiser (2014), it is plausible to speculate that for $N>10^{3}$, the connectivity may no longer be approximated by a decaying exponential.
All of the above, point to the importance of obtaining larger parcellations, providing further constraints to improve future mesoscale models of the global brain activity.
6 Conclusions
In conclusion, we have shown that an Hopfield recurrent neural network, informed by experimentally derived brain topology, recovers the scaling picture recently proposed by Deco and Kringelbach (2020), according to which the process of information exchange within the human brain shows spatially correlated patterns qualitatively similar to those displayed by turbulent flows. Our analysis confirms the initial finding by Deco and Kringelbach (2020) in near-quantitative form, predicting a very similar scaling exponent ($2/5$ instead of $1/2$) for the same value of the connectivity decay length. Given the very different nature of the two mathematical models, this provides a significant hint at the robustness of the scaling picture.
It is further observed that the scaling exponents are smaller than for the case of turbulence, indicating that the collective activity of the brain is more irregular (spikier) than homogeneous incompressible turbulence.
The scaling exponents show a steep dependence on the spatial range of neural connections, hence the specific value of the decay length has a major impact on the degree of irregularity (spikiness) of the collective patterns. Such sensitivity is also predicted to increase by increasing the size of the parcellation. This provides a strong indication towards the existence of an optimal connectivity length $\delta_{0}$ at which the brain shows non-smooth functional behavior, meaning that longer/shorter connectivity lengthscales would lead to excessive smoothness/roughness respectively. With $N=10^{3}$ parcels, this lengthscale is about $\delta_{0}\sim 5.5$ mm, well below the global size of the brain. For larger parcellations, both the transition length $\delta_{0}$ and the width of the transition region are predicted to decrease approximately like $N^{-1/3}$, which highlights the importance of assessing the optimal parcellation size.
Finally, a dilution analysis shows that, again with $N=10^{3}$, the Hopfield model brain remains functional after removing links above about six decay lengths, namely $25-30$ mm, corresponding to about one-fifth of the size of the global brain. This indicates that, in terms of connectivity decay length, the brain seems to function in a sort of intermediate “turbulent liquid-like” state, whose essential connections lie in near geometrical mean between the decay length and the global size of the brain.
7 Acknowlegments
SS wishes to acknowledge financial support from the Bertarelli Foundation during his stay at the Neurobiology Department of the Harvard Medical School. GG wishes to acknowledge financial support H2IOSC Project - Humanities and cultural Heritage Italian Open Science Cloud funded by the European Union – NextGenerationEU – NRRP M4C2 - Project code IR0000029 - CUP B63C22000730005. The authors wish to acknowledge illuminating discussions with John A. Assad, Haim Sompolinsky, and Bernardo Sabattini.